OpenStudy (anonymous):

X over 3 -6=8 sorry i didn't know how to make a fraction

OpenStudy (ash2326):

Is this your question? \[\frac{x}{3}-6=8\]

OpenStudy (anonymous):

yeaah d;

OpenStudy (ash2326):

Multiply whole equation by 3

OpenStudy (ash2326):

and post what do you get

OpenStudy (anonymous):

\[\frac{x}{3}-6(3)=8(3)\]= \[x-18=24 \] Now solve for this algebraically

OpenStudy (anonymous):

14 ?

OpenStudy (anonymous):

hmm lets see.... x-18=24 +18 +18 ____________ x=? Not quite, try following what Ive written above, and see what you get..

OpenStudy (anonymous):

i thought your subtracted 18 from both sides

OpenStudy (anonymous):

42 ^.< ?

OpenStudy (anonymous):

no you take the reciprocal of 18 to get you answer.....YEP...there you go!! in order to move -18 over to the other side you have to add it

OpenStudy (anonymous):

can you help w. this ---> Write the equation in slope-intercept form. 3x + y = 16